Find the quadratic equation which has double the roots as of equation 2x^2-5x+2=0

Grade:10

1 Answers

Nayak Roshan
7 years ago
Hi Ujjwal ,
The roots of the equation which is mentioned above are 2 and ½ 
You can find this by using factorization method 
2x2-5x+2=0
2x2-4x-x+2=0
2x (x-2)-1 (x-2)=0
THEREFORE x=2 or x=1/2.
Given to find the equation whose roots are double than that of the give equation. 
SO DOUBLE THE VALUES OF X SO OBTAINED.
THEREFORE NOW X=4 AND X=1
THEREFORE THE REQUIRED EQUATION IS x2-5x+4=0.
IF YOU ARE NOT SURE YOU CAN VERIFY THE PROBLEM BY USING THE SAME FACTORIZATION METHOD WHICH I USED ABOVE....
HOPE THIS WILL HELP YOU....